OpenStudy (anonymous):

Identify the function to match the table. A table with two columns. Left column is titled x and right column is titled y. First row shows x is negative four and y is one. Second row shows x is zero and y is two. Third row shows x is four and y is three. Fourth row shows x is eight and y is four. y= 1/4 x -1 y= 1/4 x + 2 y= 1/2 x + 5 y= 1/2 x - 3

OpenStudy (tkhunny):

Substitute the values for x and see which ones work. "First row shows x is negative four and y is one" y= 1/4 (-4) -1 = -1-1 = -2 y= 1/4 (-4) + 2 = -1+2 = 1 <== Here's your winner. y= 1/2 (-4) + 5 = -2+5 = 3 y= 1/2 (-4) - 3 = -2-3 = -5 "Second row shows x is zero and y is two" See if the same one wins again!!